Use MATLAB to plot the root locus for the system

0 0
Answer #1

Hello user,

The given euqation is 22+(0.2+k)z +6k+2 0 2+0.2c + 2 + k(z + 6) = 0 (z+6) (2+0.2z+2) So we have G(s) given by (z+6) G(s) = (2

Root Locus 8 0.91 0.84 0.74 0.56 03 0.9%8 4 8 0.99,5 20 10 2 D 0.995 4 0.98 6 0.91 0.84 0.74 0.56 03 0:955 25 8 20 0 0 5 Real

The above is obtained by command

rlocus([1,6],[1,0.2,2])

From root locus we can say that system is stable for Freq < 12.1 rad/sec

Please let me know if you have any doubts